# Novel black holes with scalar hair in the Einstein-Maxwell-Scalar Theory with positive coupling

### v1

@@ I Introduction
The advancement of astronomical detection techniques has simultaneously driven significant interest in modified gravity theories [12]. A cornerstone of GR is the no-hair theorem, which states that stationary black holes in Einstein-Maxwell theory must belong to the Kerr-Newman family, characterized solely by mass, charge, and angular momentum [13, 14, 15, 16]. However, studies have revealed that black holes in modified gravity theories can possess macroscopic external degrees of freedom. These include coupling with matter fields [17], alterations in the asymptotic structure of spacetime [18], or entirely alternative theoretical frameworks beyond Einstein–Maxwell theory [19]. The deviations of black hole solutions in modified gravity from those in pure GR serve as important tools for distinguishing extra fields and various GR extensions, as they exhibit richer physics than their bald counterparts. These deviations are predominantly manifested in the form of black hole hairs [20, 21, 22].
@@
Notably, black holes with scalar hair play a crucial role in circumventing the no-hair theorem. As the simplest matter field, scalar fields have become pivotal in various aspects of black hole physics and cosmology. They feature prominently in dark matter models [45, 46], inflation of cosmology [47, 48, 49], gravitational collapse of scalar fields and black hole formation [50, 51]. Moreover, through conformal transformations, scalar fields provide the connections between GR and modified gravity theories [52, 53, 54], which constitutes a primary motivation for investigating black hole solutions with additional scalar hair. Further details can be found in [21, 22]. Recent studies have revealed several intriguing mechanisms for black holes with scalar hair. In Kerr or charged black hole spacetimes, time-dependent bosonic fields can trigger superradiant instabilities [55, 56, 57], leading to the synchronization of the black hole [40, 58, 59] and ultimately to the formation of hairy black hole solutions. Simultaneously, studies demonstrated that a nonminimal coupling between the scalar field and the Gauss–Bonnet curvature can induce tachyonic instabilities in vacuum black holes, giving rise to spontaneous scalarization of black holes [60, 61, 62, 63]. A more straightforward approach revealed that Reissner–Nordström (RN) black holes become unstable when the scalar field is nonminimally coupled to the electromagnetic field, leading to the black hole spontaneous scalarization in the Einstein–Maxwell–scalar (EMS) theory [63, 64]. The stability of these scalarized black holes and their quasinormal modes (QNM) have been extensively analyzed in various studies [65, 66, 67, 68].
@@
In fact, in asymptotically anti-de Sitter (AdS) spacetime, holographic superconductors constitute a distinct class of hairy black holes with scalar hair. As elaborated in [69, 70], when the system temperature drops below a critical value, spontaneous symmetry breaking occurs. This manifests holographically as the scalar field acting as a vacuum expectation value begins to emerge in the boundary field theory, thereby simulating superconducting phase transitions. In the bulk, the original AdS black hole undergoes tachyonic instability, leading to the formation of a new black hole solution with scalar hair [71]. Another well-known model in holographic theories is the holographic QCD model. Recently, an improved Einstein–Maxwell–dilaton (EMD) model [72, 73, 74] has been proposed in which, via the AdS/CFT correspondence, the scalar field in the boundary theory acts as a scaling factor for the energy scale [74], successfully simulating the expected QCD phase transition and the existence of a critical endpoint. Here, the gauge field in the boundary theory corresponds to the baryon density and baryon chemical potential of the QCD model, while the black hole’s Hawking temperature and Bekenstein entropy holographically map to the temperature and entropy density, respectively. Clearly, the bulk theory necessarily contains a nontrivial scalar field to describe holographic QCD dynamics.
@@
[…]
@@ II Einstein-Maxwell-Scalar Model
[…]
@@
The event horizon is located at $r_{h}$, defined by $g(r_{h})=0$, while the asymptotic boundary corresponds to $r\rightarrow\infty$. The hawking temperature of the black hole is expressed as
@@ III Numerical Solutions
The system of field equations (13)-(16) can be numerically integrated from the event horizon to the asymptotic region using the boundary conditions (18)-(21). By employing the shooting method, we obtain numerical solutions for the field functions in the positive coupling regime ($\lambda>0$). In contrast to the negative coupling case [63, 64], we retain the first asymptotic expansion coefficient $\psi_{0}$ for the scalar field at spatial infinity. When the coupling parameter $\lambda$ remains negative, the sign of the horizon expansion coefficient $\psi_{h}^{1}$ enforces monotonic decay of the scalar field from the event horizon to the asymptotic region, allowing the shooting method to yield solutions with $\psi_{0}=0$. However, a positive coupling parameter reverses the sign of $\psi_{h}^{1}$, which results in a monotonically increasing scalar field profile. Consequently, no regular solutions exist for $\psi_{0}=0$ in this case.
@@
This observation is noteworthy because the behavior of the scalar field implies that the system does not converge to the RN black hole at asymptotic infinity. Analysis of the metric function $\chi(r)$ and the Maxwell field $A_{t}(r)$ reveals that enhanced deviations from the RN solutions accompany increasing $q$. Remarkably, as $q$ approaches zero, the system is seen to revert to the uncharged Schwarzschild solution. These observations collectively suggest that the obtained scalar-hairy solutions constitute a distinct class from both spontaneous scalarization and RN black holes in the EMS framework.
@@
[…]
@@
The behavior of the Hawking temperature as a function of $q$ for different coupling values is illustrated in the left panel of Fig. 5, with the $T-q$ relation for RN black holes represented by the black curve. It is evident that the Hawking temperature of scalar-hairy black holes is consistently lower than that of RN black holes throughout their existence domain. For larger coupling values, temperature cannot drop to zero as $q$ increases. In Fig. 2, it was observed that scalar-hairy black holes transition smoothly into RN black holes with increasing charge. However, the temperature profiles here suggest that this state transition is accompanied by a gap in temperature, although this gap diminishes with increasing coupling and eventually becomes negligible.
@@ IV Perturbation Stability
Herein, we analyze the behavior of effective potentials as shown in Fig. 6. The scalar perturbation forms a positive potential barrier near the event horizon, with the barrier height increasing both as $q$ or $\lambda$ rises. These behaviors are markedly different from that observed in spontaneous scalarization with negative coupling [64], where the potential develops negative wells near the RN limit that could trigger instabilities. The positivity of the effective potential indicates that Eq. (34) will not have a bound state, which will lead to unstable modes with $\omega^{2}<0$. Therefore, this behavior ensures the stability of our scalar-hairy black hole solutions against scalar perturbation, distinguishing it as a stable branch of hairy solutions in EMS theory that differs from the RN black hole.
@@ V Results and further discussion
In contrast to the negative coupling case where the scalar field decays with increasing radial coordinate, a defining characteristic of these scalar-hairy black holes is the monotonic growth of the scalar field along the radial direction, asymptotically approaching a constant at infinity. Similar behavior for the scalar field profile has also been observed in the EMS theory formulated in an asymptotically AdS background [75]. This phenomenon arises from the boundary conditions imposed on the scalar field near the event horizon, where the coupling parameter’s sign determines the sign of the first-derivative term in Eq. (18), governing scalar field evolution near the horizon.
@@
Numerical analysis reveals competitive interplay between scalar and electromagnetic fields for the scalar-hairy black holes. An increase in the charge $q$ causes the scalar-hairy black hole solutions to deviate further from the RN black hole, though overcharging drives the system back to the electromagnetic vacuum state. Contrary to spontaneous scalarization, where enhanced coupling enlarges the existence domain of scalarized black holes, our results show that an increase in $\lambda$ compresses the region admitting the scalar-hairy black hole state. Furthermore, a positive coupling parameter effectively weakens the interaction strength. As $\lambda\rightarrow\infty$, the system reduces to Einstein-Hilbert gravity with a massless scalar, where no stable hairy solutions exist [79]. Consequently, as shown in Fig. 4, the scalar-hairy black hole state occupies a parameter subspace within the RN boundary, whereas in the scalarization case an increased coupling can drive the scalarized branch into the $q>1$ regime. Combined with the results of Fig. 2, this indicates that the scalar-hairy black holes cannot be viewed as the endpoint of the RN black holes but might serve as a potential starting point for RN evolution in the EMS theory. Finally, our stability analysis reveals that the effective potential for scalar perturbations on the numerical background is positive, and that the associated QNMs we obtained have negative imaginary parts. These results further confirm our scalar-hairy solutions form a stable branch of hairy black holes distinct from RN family.

### v3

@@ I Introduction
The advancement of astronomical detection techniques has simultaneously driven significant interest in modified gravity theories [12]. A cornerstone of GR is the no-hair theorem, which states that stationary black holes in Einstein-Maxwell theory must belong to the Kerr-Newman family, characterized solely by mass, charge, and angular momentum [13, 14, 15, 16]. However, studies have revealed that black holes in modified gravity theories can possess macroscopic external degrees of freedom. These include coupling with matter fields [17], alterations in the asymptotic structure of spacetime [18], or entirely alternative theoretical frameworks beyond Einstein–Maxwell theory [19]. The deviations of black hole solutions in modified gravity from those in pure GR serve as important tools for distinguishing extra fields and various GR extensions, as they exhibit richer physics than their bald counterparts [20, 21]. These deviations are predominantly manifested in the form of black hole hairs [22, 23, 24].
@@
Notably, black holes with scalar hair play a crucial role in circumventing the no-hair theorem. As the simplest matter field, scalar fields have become pivotal in various aspects of black hole physics and cosmology. They feature prominently in dark matter models [47, 48], inflation of cosmology [49, 50, 51], gravitational collapse of scalar fields and black hole formation [52, 53]. Moreover, through conformal transformations, scalar fields provide the connections between GR and modified gravity theories [54, 55, 56, 57, 58], which constitutes a primary motivation for investigating black hole solutions with additional scalar hair. Further details can be found in [23, 24]. Recent studies have revealed several intriguing mechanisms for black holes with scalar hair. In Kerr or charged black hole spacetimes, time-dependent bosonic fields can trigger superradiant instabilities [59, 60, 61], leading to the synchronization of the black hole [42, 62, 63] and ultimately to the formation of hairy black hole solutions. Simultaneously, studies demonstrated that a nonminimal coupling between the scalar field and the Gauss–Bonnet curvature can induce tachyonic instabilities in vacuum black holes, giving rise to spontaneous scalarization of black holes [64, 65, 66, 67]. A more straightforward approach revealed that Reissner–Nordström (RN) black holes become unstable when the scalar field is nonminimally coupled to the electromagnetic field, leading to the black hole spontaneous scalarization in the Einstein–Maxwell–scalar (EMS) theory [67, 68]. The stability of these scalarized black holes and their quasinormal modes (QNM) have been extensively analyzed in various studies [69, 70, 71, 72].
@@
In fact, in asymptotically anti-de Sitter (AdS) spacetime, holographic superconductors constitute a distinct class of hairy black holes with scalar hair. As elaborated in [73, 74], when the system temperature drops below a critical value, spontaneous symmetry breaking occurs. This manifests holographically as the scalar field acting as a vacuum expectation value begins to emerge in the boundary field theory, thereby simulating superconducting phase transitions. In the bulk, the original AdS black hole undergoes tachyonic instability, leading to the formation of a new black hole solution with scalar hair [75]. Another well-known model in holographic theories is the holographic QCD model [76, 77, 78, 79]. Recently, an improved Einstein–Maxwell–dilaton (EMD) model [80, 81, 82] has been proposed in which, via the AdS/CFT correspondence, the scalar field in the boundary theory acts as a scaling factor for the energy scale [82], successfully simulating the expected QCD phase transition and the existence of a critical endpoint. Here, the gauge field in the boundary theory corresponds to the baryon density and baryon chemical potential of the QCD model, while the black hole’s Hawking temperature and Bekenstein entropy holographically map to the temperature and entropy density, respectively. Clearly, the bulk theory necessarily contains a nontrivial scalar field to describe holographic QCD dynamics.
@@
$S=\int dx^{4}\left(R-\frac{1}{2}\nabla^{\mu}\psi\nabla_{\mu}\psi-\frac{1}{4}f(\psi)F_{\mu\nu}^{2}-V(\psi)\right).$ (1)
@@
Opposite coupling coefficients in spontaneous scalarization have also garnered considerable research interest. Similar to neutron star scalarization, where a coupling parameter $\beta<0$ triggers instabilities in scalar perturbations, subsequent studies have shown that sufficiently large positive couplings can likewise destabilize the equilibrium configurations [84, 85]. In static black holes undergoing spontaneous scalarization, however, such opposite-sign couplings typically inhibit the development of tachyonic instabilities. By introducing angular momentum as a new degree of freedom into the theories, spin-induced spontaneous scalarization can extend the region of tachyonic instability into the regime of opposite-sign couplings, as demonstrated in ESGB gravity [86, 87] and EMS theory [88, 89]. Further details can be found in review [90]. On the other hand, beyond the framework of scalarization, the possibility of alternative mechanisms capable of generating hairy black holes remains not yet fully explored.
@@ II Einstein-Maxwell-Scalar Model
$S=\int dx^{4}\left(R-\frac{1}{2}\nabla^{\mu}\psi\nabla_{\mu}\psi-\frac{1}{4}f(\psi)F_{\mu\nu}^{2}\right),$ (2)
@@
The event horizon is located at $r_{h}$, defined by $g(r_{h})=0$, while the asymptotic boundary corresponds to $r\rightarrow\infty$. The hawking temperature of the black hole is expressed as [82]
@@ III Numerical Solutions
The system of field equations (13)-(16) can be numerically integrated from the event horizon to the asymptotic region using the boundary conditions (18)-(21). By employing the shooting method, we obtain numerical solutions for the field functions in the positive coupling regime ($\lambda>0$). It is important to note that the leading asymptotic coefficient of the scalar field, $\psi_{0}$, is retained at spatial infinity. When the coupling parameter $\lambda$ remains negative, the sign of the horizon expansion coefficient $\psi_{h}^{1}$ enforces monotonic decay of the scalar field from the event horizon to the asymptotic region [67, 68]. However, a positive coupling parameter reverses the sign of $\psi_{h}^{1}$, resulting in a monotonically increasing scalar field profile. Consequently, in this scenario, no solution exists with $\psi_{0}=0$. Nevertheless, it should be noted that asymptotically nonvanishing scalar field values are still allowed, even in scenarios involving spontaneous scalarization [91, 92].
@@
This observation is noteworthy because the behavior of the scalar field implies that the system does not converge to the RN black hole at asymptotic infinity. Analysis of the metric function $\chi(r)$ and the Maxwell field $A_{t}(r)$ reveals that enhanced deviations from the RN solutions accompany increasing $q$. Remarkably, as $q$ approaches zero, the system is seen to revert to the uncharged Schwarzschild solution. In the Schwarzschild case, the trivial scalar field solution possesses shift symmetry $\psi\rightarrow\psi+\mathcal{C}$ for any constant $\mathcal{C}$. When the Schwarzschild black hole is electrically charged, our scalar-hairy black holes can be interpreted as deformations of these shift-symmetric solutions. Generally, the RN black hole can also be viewed as the charged counterpart of the Schwarzschild solution; however, in this case, the scalar field clearly remains trivial. Spontaneous scalarization occurs due to the instability of the RN black hole, which leads to a very different scenario from our scalar-hairy solutions. These observations collectively indicate that the obtained scalar-hairy solutions constitute a distinct class from both scalarized and RN black holes in the EMS framework.
@@
To clarify the formation of scalar‑hairy black holes, we study the energy of the static solutions for scalar and electromagnetic fields, denoted by $E_{\psi}=\frac{1}{4\pi}\int_{r_{h}}^{\infty}T_{\mu\nu}^{\psi}n^{\mu}n^{\nu}dV$ and $E_{\text{M}}=\frac{1}{4\pi}\int_{r_{h}}^{\infty}T_{\mu\nu}^{\text{M}}n^{\mu}n^{\nu}dV$ on a normal vector of the constant time slice [93]. Fig. 3 displays the relations between the field energy and the reduce charge $q$ for different coupling, where the upper panel corresponds to spontaneous scalarization ($\lambda<0$) and the lower panel represents the scalar-hairy solutions ($\lambda>0$). During the charging process, both the scalar and maxwell fields accumulate energy, with the scalar field consistently storing less energy than its electromagnetic counterpart. However, in the spontaneous scalarization case, one sees a nonmonotonic behavior in $E_{\text{M}}$: in the small-$q$ region, the electromagnetic field energy initially decreases before beginning to increase, indicating a temporary loss of small amounts of energy in the electromagnetic field at the onset of spontaneous scalarization. Notably, increasing coupling strength in this case induces energy reduction in both scalar and Maxwell fields, whereas scalar-hairy black holes exhibit the opposite trend with enhanced energy storage at stronger coupling. This intriguing observation of opposite energy transfer directions suggests that spontaneous scalarization and scalar-hairy black holes are driven by distinct mechanisms. The spontaneous scalarization onset likely originates from tachyonic instability, while in positively coupled scalar-hairy black holes where tachyonic instability is absent, nonlinear coupling effects potentially dominate the energy dynamics.
@@
The behavior of the Hawking temperature as a function of $q$ for different coupling values is illustrated in the left panel of Fig. 5, with the TqT-q relation for RN black holes represented by the black curve. It is evident that the Hawking temperature of scalar-hairy black holes is consistently lower than that of RN black holes throughout their existence domain. For larger coupling values, temperature cannot drop to zero as $q$ increases. In Fig. 2, it was observed that scalar-hairy black holes transition smoothly into RN black holes with increasing charge. However, the temperature profiles here suggest that this state transition is accompanied by a gap in temperature, although this gap diminishes with increasing coupling and eventually becomes negligible.
@@ IV Perturbation Stability
Herein, we analyze the behavior of effective potentials as shown in Fig. 6. The scalar perturbation forms a positive potential barrier near the event horizon, with the barrier height increasing both as $q$ or $\lambda$ rises. These behaviors are markedly different from that observed in spontaneous scalarization with negative coupling [68], where the potential develops negative wells near the RN limit that could trigger instabilities. The positivity of the effective potential indicates that Eq. (34) will not have a bound state, which will lead to unstable modes with $\omega^{2}<0$. Therefore, this behavior ensures the stability of our scalar-hairy black hole solutions against scalar perturbation.
@@ V Results and further discussion
In contrast to the negative coupling case where the scalar field decays with increasing radial coordinate, a defining characteristic of these scalar-hairy black holes is the monotonic growth of the scalar field along the radial direction, asymptotically approaching a constant at infinity. Similar behavior for the scalar field profile has also been observed in the EMS theory formulated in an asymptotically AdS background [83]. This phenomenon arises from the boundary conditions imposed on the scalar field near the event horizon, where the coupling parameter’s sign determines the sign of the first-derivative term in Eq. (18), governing scalar field evolution near the horizon. Furthermore, from another perspective, the scalar-hairy black hole solutions can be regarded as the charged counterparts of the Schwarzschild black hole. Although the RN black hole also represents a charged Schwarzschild solution, it does not support scalar hair in the EMS theory, which distinguishes it from our scalar-hairy black holes.
@@
Numerical analysis reveals competitive interplay between scalar and electromagnetic fields for the scalar-hairy black holes. An increase in the charge $q$ causes the scalar-hairy black hole solutions to deviate further from the RN black hole, though overcharging drives the system back to the electromagnetic vacuum state. Contrary to spontaneous scalarization, where enhanced coupling enlarges the existence domain of scalarized black holes, our results show that an increase in $\lambda$ compresses the region admitting the scalar-hairy black hole state. Furthermore, a positive coupling parameter effectively weakens the interaction strength. As $\lambda\rightarrow\infty$, the system reduces to Einstein-Hilbert gravity with a massless scalar, where no stable hairy solutions exist [96]. Consequently, as shown in Fig. 4, the scalar-hairy black hole state occupies a parameter subspace within the RN boundary, whereas in the scalarization case an increased coupling can drive the scalarized branch into the $q>1$ regime. Combined with the results of Fig. 2, this indicates that the scalar-hairy black holes cannot be viewed as the endpoint of the RN black holes but might serve as a potential starting point for RN evolution in the EMS theory. Finally, our stability analysis reveals that the effective potential for scalar perturbations on the numerical background is positive, and that the associated QNMs we obtained have negative imaginary parts. These results confirm that our scalar-hairy solutions as a new branch of hairy black holes distinct from RN family, remain linearly stable at least against scalar perturbations. A comprehensive stability assessment of these scalar‑hairy black holes requires further examination under either gravitational or electromagnetic perturbations [71, 97, 98]. Furthermore, a more detailed discussion of the thermodynamic examination and interrelations of scalar-hairy and RN black holes in the vicinity of the existence line will further elucidate their properties and the potential black hole phase transitions.